# Multimodal AI predicts clinical outcomes of drug combinations from preclinical data

### v1

@@ Results
We used Madrigal to examine a common mechanism of DDIs—shared transport mechanisms among drugs [57, 58]. Membrane transporters play a crucial role in drug absorption, distribution, and elimination, and their shared use by multiple drugs can lead to significant DDIs. A comprehensive study by the International Transporter Consortium [59] revealed that approximately 75% of the top 200 prescribed drugs are substrates for at least one transporter, and many of these drugs interact with multiple transporters, creating a high potential for transporters to play a role in the governance of drug absorption and disposition and in mediating clinical DDI.
Transporter-mediated DDIs [57], such as doxycycline’s interactions with digoxin, warfarin, tacrolimus, and levetiracetam, are effectively captured by Madrigal (Fig. 3d). Specifically, Madrigal assigned a very high normalized rank (prediction score ranked among all DrugBank drugs then normalized to [0,1], see Methods. 5.1), indicating a less favorable safety profile, to the doxycycline-tacrolimus pair and a moderately high rank to doxycycline-levetiracetam, despite neither pair being present in Madrigal’s dataset. Drugs that share transporters, enzymes, or carriers showed a significantly higher tendency for interaction in Madrigal’s predictions (Fig. 3e). For drugs that share specific transporters, which the regulatory guidelines recommend for evaluation due to organ-specific safety risks[58], Madrigal also highlighted transporter-related safety events (Fig. 3f). This accuracy underscores Madrigal’s potential as a tool for preemptively identifying risks for drug interactions early in drug development both from a clinical pharmacology perspective and in combination drug development.
Madrigal predicts safety profiles of oncology drug combinations.
Balancing efficacy and safety in combination cancer therapies is challenging due to the severe toxicity profiles of treatments such as chemotherapy, where safety testing often restricts which combinations progress to efficacy studies [60]. We investigate Madrigal for combinations that include PARPi [^1] Included in the analysis were PARPi’s that are FDA-approved or in late-phase clinical trials and are sufficiently represented in the dataset (with at least five different drug combinations and data for at least two outcomes)., a class of targeted therapies used to treat breast and ovarian cancer [61].
We first examine whether known neutropenia and hematological effects of PARPi combinations are captured by Madrigal (as shown in Fig. 4a). The olaparib + paclitaxel combination, with a predicted safety rank exceeding 750, was associated with higher rates of grade 3 and 4 hematological adverse effects, such as neutropenia, in gastric cancer patients [62]. In contrast, the combination of olaparib + abiraterone, which had a predicted safety rank of 300, was studied in metastatic castration-resistant prostate cancer (mCRPC) [63]. The latter combination, administered in conjunction with prednisone or prednisolone, has been approved in the US for adult patients with deleterious or suspected deleterious BRCA-mutated mCRPC (BRCAm) [64], and in the EU for adult male patients with mCRPC in whom chemotherapy is not clinically indicated [65]. The adverse effects of drug combinations, including hematologic toxicities, can be explained by the pharmacological properties of the drugs in addition to the symptoms or etiology of the disease [66]. Hematological toxicities, such as neutropenia, are largely attributable to drug-induced myelosuppression, a common property of chemotherapy agents. These findings indicate that the safety classifications predicted by Madrigal capture broad hematological adverse events of drug combinations that are generalized between types of cancer and patient populations.
The combination of olaparib + paclitaxel is predicted to have a higher risk of hematological and neutropenia-related adverse events than olaparib + ceralasertib, which is a combination that is being clinically investigated with incomplete safety characterization (Fig. 4a shows median-aggregated scores across neutropenia and hematological effects, Supplementary Fig. S5a shows mean-aggregated scores). In plasmaMATCH [67], CAPRI [68, 69], and VIOLETTE [70] studies (up to Phase II), the combination of olaparib + ceralasertib showed a manageable safety profile, with adverse events consistent with those expected from each drug individually and no new significant safety concerns. The combination of olaparib + paclitaxel, however, evaluated in the Phase III GOLD trial [62], resulted in grade 3 or worse neutropenia adverse events in 30% of patients with advanced gastric cancer. The olaparib + abiraterone combination has received FDA approval [64], with 5% of mCRPC patients experiencing grade 3 or worse neutropenia in the Phase III PROpel trial [63].
Of note, our dataset does not include annotated outcome data for the olaparib + abiraterone and olaparib + ceralasertib combinations, while the olaparib + paclitaxel combination is only annotated with an increase in metabolic activity. The hematological annotations for olaparib in the dataset include an increased risk of bleeding, methemoglobinemia, neutropenia, and thrombocytopenia, and these effects are documented for combinations involving only 12 other drugs. Other PARP inhibitors, such as rucaparib and talazoparib, have even fewer annotations related to hematological outcomes in the dataset. Combinations of these drugs and local anesthetics are associated with an increased risk of methemoglobinemia, with no other documented hematological effects, and methemoglobinemia is a rare adverse effect associated with local anesthetics [71, 72]. In addition to the prediction of increased risk of neutropenia (0.998, 8/141) for olaparib + paclitaxel as presented above, metabolism, excretion, myelosuppression, gastrointestinal ulcers, cardiotoxicity, and liver damage are among the top 10 adverse events predicted by Madrigal. The prediction of liver damage aligns with the complications observed in the GOLD study. In the (olaparib + paclitaxel) treatment arm, liver damage was reported as a treatment-related adverse event and was implicated in the death of one patient [62].
Anemia is a well-documented adverse event associated with PARPi’s and can be clinically managed [64]. Since the dataset of Madrigal does not include any associations between PARPi and anemia, this adverse effect is used as a positive control to evaluate Madrigal to detect effects that occur in patients, but are not included in the dataset. We prompted Madrigal-LLM with a description of anemia (Fig. 7f). Madrigal predicted significantly higher scores for anemia in olaparib + abiraterone combination compared to random drug combinations ($p$-value $<0.05$, two-sided Mann-Whitney U).
Next, we investigate organ-specific effects of PARPi combined with other anti-cancer drugs (Methods Sec. 5.3). For each PARPi, we use Madrigal to make predictions for its combination with all other anti-cancer drugs we curated across hematological, renal, and hepatic outcomes (Fig. 4b, Methods Sec. 5.3). Hematological adverse events are well-documented side effects of PARPi, both when used as monotherapies and in combination therapies [64, 73, 74]. Among these, ”neutropenia, increase” is consistently ranked as a highly predicted effect for all PARPi combinations (olaparib: 7/141; talazoparib: 9/141; rucaparib: 3/141). Systematic review of hematological adverse events [75] shows minor differences between olaparib and rucaparib. Renal adverse events, such as increased blood creatinine and acute kidney injury, have been reported in the TRITON2 trial of rucaparib [74] and in eight clinical trials of olaparib [64], including the PROpel trial investigating the olaparib + abiraterone combination [63]. Four of these eight trials reported that over 25% of participants experienced an increase in serum creatinine. Although Madrigal’s dataset does not include these specific renal events, Madrigal predicts high scores for renal-related adverse events across PARPi combinations, demonstrating its ability to generalize beyond the training data. These include changes in excretion (1/141 in all three PARPi’s), serum levels (talazoparib: 4/141; rucaparib: 14/141), serum levels excretion rate (olaparib: 5/141; talazoparib: 6/141; rucaparib: 5/141) and serum levels of active metabolites (olaparib: 14/141; talazoparib: 17/141; rucaparib: 13/141), indicating a potential risk of renal safety events. The EMBRACA trial [76] of talazoparib in patients with advanced breast cancer and germline mutations in BRCA1/2 reported one drug-related death in the talazoparib group due to veno-occlusive liver disease. Increased aminotransferase levels were identified as common adverse reactions (occurring in more than 20% patients) in the EMBRACA trial (aspartate and alanine aminotransferases) and in the GOLD trial of olaparib combined with paclitaxel, where elevations of alanine aminotransferase were reported in more than 10% of patients [62].
The renal and hepatic effects described above are also associated with transporter inhibition, a characteristic of olaparib and rucaparib [64, 74]. Differences in predicted renal and liver outcomes for talazoparib compared to rucaparib and olaparib (aggregated scores of 0.937 vs. 0.955 and 0.947 for renal effects; 0.939 vs. 0.958 and 0.957 for liver effects, Figure 4b) align with in vitro findings indicating distinct inhibitory profiles of these PARP inhibitors on enzymes and transporters. Specifically, olaparib and rucaparib act as inhibitors and inducers of multiple enzymes and as inhibitors of various transporters [64, 74], whereas talazoparib has minimal impact on enzymes and transporters [73]. These differences raise the possibility of olaparib- and rucaparib- (but not talazoparib-) specific drug interactions with enzyme- and transporter-related mechanisms, which have also been suggested by pharmacokinetic studies [77].
Drug combinations approved or under clinical investigation have undergone a thorough safety investigation. As such, we next seek to examine whether drug combinations involving PARPi (“PARPi combination”) that are being clinically investigated or approved are predicted by Madrigal as having more favorable safety profiles. PARPi combinations in which the other drug is any cancer drug are treated as a negative control. We found that those clinically investigating or approved PARPi combinations, curated from [78, 64, 73], are predicted by Madrigal to have a more favorable safety profile than PARPi combinations (with any other cancer drug) in endocrine, heart, kidney, and liver effects, and comparable in blood and gastrointestinal effects (Fig. 4c). In all organs except the liver, these combinations are predicted to have a more favorable safety profile compared to clinically investigated drug combinations, including those actively investigated in 2024, those that have failed clinical trials or those that have been withdrawn [79]. Drug combinations already used in patients and listed in the US FDA Orange Book are predicted to exhibit the most favorable safety profiles across all organs, with the exception of the heart, which is used as a positive control.
We visualize the normalized ranks of all outcomes for each PARPi combination currently under clinical investigation, including the (olaparib + paclitaxel) combination to serve as a reference for a combination with known safety concerns. All PARPi combinations under investigation are predicted by Madrigal to have a more favorable safety profile compared to the control (Fig. 4d, from left to right with increasing average of the top five highest normalized ranks). Based on this average, we observe that PARPi combinations that have progressed to advanced stages of clinical development, as well as those that are approved, are generally predicted to have more favorable safety profiles than those in earlier phases. Since early-stage trials, particularly phase I, primarily focus on assessing candidate safety, these combinations often have less well-established safety profiles compared to those in advanced phases or approved for use. The predictions made by Madrigal align with this trend, as the variations in safety profiles among phase I combinations suggest differences in their suitability for clinical development.
Polypharmacy management in chronic metabolic disorders using Madrigal.
The management of chronic metabolic disorders often involves complex polypharmacy due to comorbid diseases that require the concurrent use of multiple drugs. The incomplete mechanistic understanding of many therapies, coupled with the intricate and overlapping pathophysiology of these disorders, underscores the importance of combination therapies [80] This complexity is particularly evident in conditions such as type II diabetes (T2D) and MASH, both of which are increasingly prevalent [81, 82] and characterized by multifaceted pathophysiology [83, 3, 84, 85]. MASH is among the most frequently cooccurring diseases in patients with T2D [82], with a global prevalence that increased from 5-7%, to more than seven times higher, 37%, among people with T2D patients [86], and occurs in up to 70% of patients with T2D in specific populations analyzed [87]. Complicating the situation even more, the FDA approved the first drug, resmetirom, for MASH only in 2024 [88]. Given the pressing need for more effective treatment protocols, combination therapies are being considered for T2D and MASH. We use Madrigal to investigate drug combinations in three related areas: (1) combinations for T2D and heart failure (T2D + HF), a better understood comorbidity; (2) combinations for T2D and MASH (T2D + MASH), a complex and emerging area of focus; and (3) combinations for MASH alone, where treatment options are particularly limited. To support our analysis, the T2D medications are sourced from DrugCentral [89] through PrimeKG knowledge graph [90], the Orange Book of the US FDA, and supplemented with mechanism of action data from UCSF Diabetes Teaching Center [91], Mayo Clinic [92], and Cleveland Clinic [93], and DrugBank [46]. We ensure that all medications, if not approved in the US, are marketed in Europe or Japan. Due to our focus on small molecule drugs, insulin and its analogs are excluded from this analysis.
@@
First, we seek to validate that the predicted safety profiles of drug combinations for T2D and its comorbidities align with clinical knowledge. In the case of heart failure comorbidity, consistent with clinical observations [94], combinations involving rosiglitazone with heart failure medications are predicted to have worse safety profiles in multiple effects related to myocardial infarction or stroke compared to combinations involving pioglitazone (Fig. 5a). Madrigal’s predictions also reflect the known risks of hyperkalemia associated with renin-angiotensin-aldosterone system (RAAS) inhibitors, including aldosterone antagonists, angiotensin II receptor blockers (ARBs), angiotensin receptor-neprilysin inhibitors (ARNi’s), and ACE inhibitors (ACEi’s)[95], as well as the observation that SGLT2 inhibitors may reduce the risk[96]. Combinations of these drugs with sodium zirconium cyclosilicate (SZC), a drug used to treat hyperkalemia, show significantly improved safety profiles compared to other combinations, which aligns with recent clinical practice [97, 98] and trial design [99]. Candesartan, which has a known risk of hyperkalemia [100], was used as a control (Fig. 5b). For renal effects, our model predicts that combinations of T2D medications with SGLT2 inhibitors have more favorable renal safety profiles compared to combinations with diuretics, consistent with reported clinical observations [101, 102] (Supplementary Fig. S6a).
Next, we focus on (T2D + MASH) combinations in light of the unmet need for MASH therapies. We evaluate the safety profiles of all MASH drugs or candidates currently approved, under clinical investigation, or being used off-label [103], when combined with any T2D drugs that do not share the same MoA (“rational combination”, Fig. 5c,g). The complete predicted ranks are provided in Supplementary Fig. S6b. We also calculate the normalized rank for each MASH drug or candidate in a minimax manner (Methods Sec. 5.4). Madrigal ranks resmetirom, the first and only FDA-approved drug for metabolic dysfunction-associated steatohepatitis (MASH) as having the second most favorable safety profile. The predicted candidate with the most favorable safety profile, elafibranor, has demonstrated a consistent safety profile in various trials, including those for primary biliary cholangitis (PBC)[104]. However, it should be noted that in the RESOLVE-IT Phase III trial (NCT02704403), elafibranor did not show a statistically significant effect on resolving MASH without worsening fibrosis, although its safety and tolerability were consistent with previous studies. This suggests that while single-drug elafibranor efficacy in MASH might be insufficient, its safety profile may still hold potential for further exploration in combination therapies for MASH. Many of the candidates predicted to have more favorable safety profiles have safety records without serious adverse effects for effective doses in clinical trials [104, 105, 106]. While the model predicts firsocostat to have relatively favorable safety profile as well, real-world evidence suggests otherwise, particularly its association with hypertriglyceridemia [107], which is not captured in Madrigal dataset due to a lack of relevant effects annotations, and that firsocostat itself has no annotations of any kind. We then examine the clinical trial phases of MASH drugs or candidates and observe that when combined with T2D drugs, candidates in phase I—where the primary focus is on assessing safety—tend to exhibit less favorable safety profiles compared to those in later phases (II to IV), as expected (Fig. 5d,e). Additionally, we observed that candidates with different MoAs demonstrate varying safety profiles when combined with T2D drugs (Fig. 5f), highlighting the influence of MoA on drug safety in combination therapies.
In addition to developing monotherapies for MASH, multiple trials have explored combination therapies to understand the mechanisms underlying the safety and efficacy of drug combinations. The rationales behind MASH combination therapies broadly fall into two categories [3]: (1) Two drugs target independent pathways or two components of one pathway, leading to improved efficacy through independent action or synergy [1], especially given the complexity of MASH (enhanced efficacy); (2) A drug addresses MASH, while a secondary drug mitigates the adverse effects associated with the primary therapy (enhanced safety) (Fig. 5g). We annotate drug combinations currently under clinical investigation for MASH [108, 85, 84, 3] into these two categories. While the annotation of enhanced safety or efficacy was based on literature curation, these categories are not mutually exclusive—some combinations may confer both safety and efficacy benefits. We then ranked the safety of these combinations based on the average of the top five normalized ranks among all possible pairwise combinations of MASH drugs or candidates within the same MoA class (“background combinations”), representing relative safety among possible alternative combinations that share similar combination rationales. We observe that 3 out of 5 combinations curated for enhanced safety and 5 out of 11 combinations curated for enhanced efficacy ranked relatively safe, which means that they were ranked first among fewer than five background combinations or ranked first or second in no fewer than five background combinations, respectively (Fig. 5h).
Predicting individualized drug combination efficacy in cancer patients with Madrigal.
The clinical effectiveness of cancer combination therapies can depend on the tumor genomic environment, highlighting the need to identify genomic signatures that influence therapy response [1, 109, 60]. Building on Madrigal’s ability to predict outcomes, we extended Madrigal to predict the efficacy of drug combinations in cancer patients. We adapted Madrigal to predict the efficacy and synergy of drug combinations in ex vivo cancer models (Fig. 6b). To validate Madrigal for personalized drug response predictions, we used drug response data from patient-derived xenografts (PDXE [41]) and primary cancer cells (BeatAML [39, 40]) (Fig. 6a, Supplementary Fig. S7b). This approach can optimize combination therapies tailored to the unique needs of each patient, factoring in safety and efficacy, two key pillars of success in drug discovery and development.
@@
In the BeatAML dataset, we sampled patients to hold out for evaluation based on gene expression profiles, somatic mutation data, and clinical information with less than 10% missing data (Methods Sec. 5.6). For each drug combination, Madrigal’s drug embeddings were fed into a bilinear decoder to generate a fused embedding, which was then concatenated with dimension-reduced gene expression and clinical attribute vectors before being inputted into a multi-layer perceptron. The classification target was defined as synergistic if the “combination ratio” was less than 1 [40] (Methods Sec. 1.9). Multimodal Madrigal combined with patient genomic profiles outperformed models without genomic input, showing the importance of patient data for strong performance (Fig. 6c). However, patient information alone was insufficient to achieve meaningful predictions without combined drug embeddings.
In the PDX dataset, we incorporated gene expression and mutation data into Madrigal, testing Madrigal’s performance by holding out each drug combination. For this dataset, we used the element-wise maximum of the two drug embeddings from Madrigal, concatenated with dimension-reduced gene expression data, as input to a random forest regressor. We trained two models with different regression targets: treatment response (“BestAvgResponse” [41]) and progression-free survival (PFS, “TimeToDouble” [41]). When predicting response, Madrigal showed statistically significant correlations between predicted and actual responses for 8 of 11 held-out drug combinations ($p$-value $<0.05$, Kendall’s $\tau$), with an average correlation of $\tau=0.466$ (Fig. 6d).
We stratified patients into two response categories based on predicted response using the mRECIST criterion [41]: patients with a predicted response below -20 were classified as responders and others as non-responders. For drug combinations with sufficient predicted patient responders and non-responders ($\geq 5$), Kaplan-Meier survival estimates showed statistically significant differences between these subgroups based on ground truth PFS (Fig. 6e,f, Supplementary Fig. S7c). When predicting PFS, Madrigal showed a significant correlation with outcomes for 7 of 11 drug combinations in the dataset ($p$-value $<0.05$, Kendall’s $\tau$), with an average correlation of 0.489 (Supplementary Fig. S7d). Using the same response-based stratification approach, we confirmed that the PFS predictions aligned with the observed patient response outcome, further validating Madrigal’s performance (Fig. 6g,h, Supplementary Fig. S7e). Madrigal integrates tumor genomic profiles to predict the outcomes of anti-cancer drug combinations. Using these profiles, Madrigal identifies patients who are more likely to benefit from specific combination therapies, allowing a more personalized approach to cancer treatment.
Enhancing Madrigal with an LLM for new drug combination descriptions.
We integrated Madrigal with a text-based LLM to address drug combination effects and outcomes beyond those encountered in Madrigal’s training data or codified in existing medical vocabularies. Existing models are limited to controlled vocabularies, restricting their ability to predict a wider range of outcomes, highlighting the need for a flexible model that maintains predictive precision in diverse outcomes [110]. LLM-only predictors use drug names [111, 112] in input prompts, which limits their ability to be used with new compounds in development that lack designated names assigned post-approval [113, 114] or compounds with limited information in literature and thus poor representation in LLM’s knowledge base [115].
To address this, we designed a modeling approach that combines Madrigal with an LLM, resulting in Madrigal-LLM. This model leverages text descriptions of drug outcomes from DrugBank alongside drug embeddings from contrastive pretrained Madrigal (without finetuning on drug combination datasets) as input to predict the likelihood of a drug pair linked to a specific outcome (Fig. 7a). Using a text-based LLM, these descriptions are converted into outcome embeddings. Placeholder tokens replace drug names in these text inputs, ensuring the LLM does not rely on specific drug names or leak information from its internal knowledge base based on drug names mentioned in the user prompt [115]. To test Madrigal-LLM’s generalizability, we held out outcomes so that those in the test set were not encountered by Madrigal-LLM during training. To further enrich the diversity of text input, we used GPT-4 to generate augmented descriptions from the perspectives of experts in medicine, pharmacology, and toxicology (Fig. 7b). This strategy exposed Madrigal-LLM to a spectrum of text variations in outcomes. During training, one of the ten possible descriptions (the original plus nine expert-augmented descriptions) was randomly sampled for each drug combination, allowing the model to learn from various linguistic inputs at each training step. Performance is assessed across all text descriptions for each test outcome. We included a reference score, the likelihood-weighted prior (Methods Sec. 4), which estimates the probability of a drug combination outcome in the test set based on its occurrence in the training set, adjusted for semantic similarity to known outcomes.
We evaluated whether augmented text descriptions can enhance Madrigal-LLM performance trained solely on original (non-augmented) outcome descriptions. Across outcomes (13/16), models trained with augmented descriptions performed better than those trained without augmentation (Fig. 7c, with three random initializations for each model). Both models trained on original and augmented descriptions performed better than baseline in most outcomes (10/16 and 12/16, respectively). We further tested Madrigal-LLM trained on augmented text against the original and paraphrased descriptions in the test set (Fig.7d). In most outcomes (13/16), at least one expert-generated description improved performance over the original description, indicating the added value of expert-driven text diversity. Finally, we analyzed Madrigal-LLM’s predictions by measuring concordance between Madrigal-LLM and Madrigal. Using Spearman’s rank correlation, we compared predicted scores from both models for outcomes unseen by Madrigal-LLM (Fig. 7e, showing only correlations with $p$-value $<$ 0.05 due to limited sample size in some outcomes). For most outcomes, strong positive correlations are observed, indicating that Madrigal-LLM effectively captures outcome patterns similar to those learned by Madrigal, even without explicit training on those drug combination outcomes.

### v2

@@ Results
We next examine a common mechanism of DDIs: shared transport mechanisms [58, 59]. Membrane transporters influence absorption, distribution, and elimination, and shared transporter use can produce clinically meaningful interactions. Analysis by the International Transporter Consortium reported that about 75% of the top 200 prescribed drugs are substrates of at least one transporter, with many engaging multiple transporters and therefore at higher potential for transporter-mediated DDI [60].
Madrigal captures transporter-mediated DDIs, exemplified by doxycycline’s interactions with digoxin, warfarin, tacrolimus, and levetiracetam (Fig. 3d). Madrigal assigns a high normalized rank (prediction score ranked among all drugs and normalized to [0,1]; Methods Sec. 4.1) to the doxycycline + tacrolimus pair and a moderately high rank to doxycycline + levetiracetam, despite neither pair appearing in Madrigal’s training data.
Drugs that share transporters, enzymes, or carriers show a significantly higher tendency for interaction in Madrigal’s predictions (Fig. 3e). For drugs sharing specific transporters that regulatory guidance prioritizes due to organ-specific safety risks [59], Madrigal highlights corresponding transporter-related safety events (Fig. 3f). These results suggest that Madrigal can help prioritize potential transporter-mediated risks for follow-up testing and clinical risk management.
Alignment with clinical trial safety.
Controlled trials that compare combinations head-to-head provide a rigorous benchmark for combination toxicity. We identify 35 advanced-stage trials since 2000 that tested multiple small-molecule combinations under comparable conditions (Methods Sec. 1.8; Supplementary Table S10). For each trial, we select adverse events (AEs) with significantly different incidences between arms (Methods Sec. 4.3). Across AEs with at least five arm comparisons (alopecia, anemia, hypoglycemia, and neutropenia), Madrigal’s ordering of arm-pair risk agrees with trial outcomes in 7/8, 4/5, 6/6, and 7/9 arm pairs, respectively (Fig. 4a; Supplementary Tables S11-S14, 19/35 trials with significantly different incidences between arms for at least one of neutropenia, hypoglycemia, anemia, and alopecia). Agreement is defined as the trial arm with a more favorable safety profile also receiving a lower Madrigal score (Methods Sec. 4.3).
We next test whether Madrigal differentiates combinations that have progressed into the clinic. Madrigal recapitulates known hematologic toxicities of poly(ADP-ribose) polymerase inhibitor (PARPi) combinations, including the higher rates of grade 3-4 hematologic AEs, including neutropenia, reported for (olaparib + paclitaxel) in gastric cancer [61], which has so far not been approved (Supplementary Fig. S5a-c; Supplementary Note S9).
PARPi combinations under investigation or approval [62, 63, 64] are predicted to have more favorable safety profiles than pairing PARPi with cancer drugs for endocrine, kidney, heart, and liver effects, and to be comparable for blood and gastrointestinal effects (Fig. 4c; Supplementary Note S10). In all organs except liver, these PARPi combinations are also predicted to be safer than clinically investigated oncology combinations, including those active in 2024, failed, or withdrawn [65]. Combinations already used in patients (US FDA Orange Book) are predicted to be safest overall across organs, with heart as second safest. We visualize normalized ranks across outcomes for each PARPi combination under clinical investigation. PARPi combinations that have advanced further clinically or are approved generally receive more favorable safety predictions than those in earlier phases (Supplementary Fig. S5d; ordered left to right by increasing average of the top five highest normalized ranks).
Applying Madrigal to T2D and MASH polypharmacy.
The management of chronic metabolic disorders often requires complex polypharmacy due to multimorbidity and multifactorial pathophysiology [66]. This is pronounced in type 2 diabetes (T2D) and metabolic dysfunction-associated steatohepatitis (MASH), which are increasingly prevalent [67, 68] and heterogeneous in mechanism [69, 3, 70, 71]. MASH frequently co-occurs with T2D [68], with global prevalence estimates rising from 5-7% in the general population to 37% among people with T2D [72]. The first MASH therapy, resmetirom, was approved in 2024 [73]. We use Madrigal to examine combinations in three settings: (1) T2D and heart failure, a well-characterized comorbidity; (2) T2D and MASH, an emerging area; and (3) MASH combination therapy, where options are limited.
@@
T2D and heart failure. Madrigal’s safety rankings align with clinical knowledge. Combinations involving rosiglitazone are predicted to have less favorable cardiovascular profiles than those with pioglitazone, consistent with reports of myocardial infarction and stroke risk [74] (Fig. 5a; Supplementary Table S4, S6). When pairing heart-failure therapies with glucose-lowering drugs, Madrigal reflects the hyperkalemia risk associated with renin-angiotensin-aldosterone system inhibitors [75] and the mitigating association of SGLT2 inhibitors [76]. Combinations including sodium zirconium cyclosilicate, used to treat hyperkalemia, are predicted to have significantly improved safety relative to other pairings, consistent with recent clinical practice and trial design [77, 78, 79]; candesartan, with known hyperkalemia risk [80], serves as a control (Fig. 5b). For renal effects, combinations with SGLT2 inhibitors are predicted to have more favorable renal safety than those with diuretics, in line with clinical observations [81, 82] (Supplementary Fig. S6a).
T2D and MASH. We evaluate safety profiles when MASH drugs or candidates (approved, in trials, or used off-label [83]) are combined with T2D drugs of different mechanisms (Fig. 5c,g; Supplementary Fig. S6b; Supplementary Table S5; Methods Sec. 4.4). Madrigal ranks resmetirom, the first FDA-approved MASH drug, as the second most favorable. The top-ranked candidate, elafibranor, has shown a consistent safety profile across trials, including in primary biliary cholangitis [84]; although the RESOLVE-IT Phase III trial (NCT02704403) did not meet its primary MASH efficacy endpoint, safety and tolerability were consistent with prior studies, suggesting a potential role for elafibranor within combination regimens. Some risks are not captured by our label sets: tropifexor’s dose-related pruritus in Phase II [85] and firsocostat-associated hypertriglyceridemia [86] are not annotated in the Madrigal DrugBank-derived outcomes, which may lead to underestimation in our predictions. Across candidates, combinations involving Phase I candidates, where clinical programs focus on safety, tend to rank less favorably than those in later phases (Fig. 5d,e), and safety varies by mechanism of action when paired with T2D drugs (Fig. 5f).
MASH combination strategies. MASH combinations are typically motivated by either improved efficacy (targeting independent pathways or multiple nodes of one pathway) or improved safety (a second agent mitigates the first agent’s adverse effects) [3] (Fig. 5g). We annotate combinations under clinical investigation [87, 71, 70, 3] and rank their safety using the average of the top five normalized ranks across outcomes. Each regimen is compared against background combinations formed from other pairings between drug pairs with the same respective mechanisms of action (Methods Sec. 4.4). By this criterion, 3/5 combinations curated for enhanced safety and 5/11 curated for enhanced efficacy rank as relatively safe (ranked first among fewer than five background combinations, or first/second among five or more; Fig. 5h).
Personalized ex vivo efficacy prediction with Madrigal.
We test whether Madrigal can support individualized prioritization of cancer drug combinations [88, 89] by pairing its drug embeddings with patient molecular profiles in two ex vivo systems: primary cancer cells from BeatAML and patient-derived xenografts from PDXE (Fig. 6a; Supplementary Fig. S7a).
@@
BeatAML. For each drug combination, Madrigal’s two drug embeddings are fused with a bilinear decoder. The fused embedding is concatenated with dimensionality-reduced gene expression and clinical attributes, then passed to a multi-layer perceptron (Methods Sec. 1.9). The classification target is synergy defined by a drug combination ratio less than 1 [40] (Supplementary Fig. S7c). Models that combine Madrigal with patient genomic profiles outperform models without genomics when patients are held out (Fig. 6b) and when drugs are held out (Supplementary Fig. S7b), indicating the predictive value of patient information. Patient features alone are insufficient without the combined drug embeddings.
PDXE. We integrate gene expression and mutation data with Madrigal and evaluate performance by holding out each drug combination. The element-wise maximum of the two Madrigal drug embeddings is concatenated with dimensionality-reduced gene expression and used as input to a random forest regressor. We train two predictors: treatment response (BestAvgResponse) and progression-free survival (PFS, TimeToDouble) [41]. For response, Madrigal shows significant correlations between predicted and observed responses in 8 of 11 held-out combinations (Kendall’s $\tau$, $p<0.05$), with an average $\tau=0.465$ (Fig. 6c). We stratify patients using an mRECIST-based threshold (predicted response below 20-20 as responders, otherwise non-responders) [41]. For combinations with at least five predicted responders and five predicted non-responders, Kaplan-Meier estimates show significant differences in ground-truth PFS between strata (Fig. 6e,f; Supplementary Fig. S7d). For PFS prediction, Madrigal again correlates with observed survival in 7 of 11 combinations ($p<0.05$, average $\tau=0.489$; Supplementary Fig. S7e). Stratifying by predicted response reproduces differences in observed PFS (Fig. 6d; Supplementary Fig. S7f). These results show that Madrigal can transfer from multimodal preclinical drug information to patient-level ex vivo efficacy when combined with genomic context.
Real-world, personalized safety prediction in EHR and oncology cohorts.
Having established that Madrigal supports individualized combination response in ex vivo models, we next evaluate clinical safety prediction in patients (Fig. 6g,h).
Longitudinal EHR cohort. We integrate Madrigal into TransformerEHR [90] and fine-tune on longitudinal health records from the EHRSHOT Stanford Medicine cohort (n=6,739) [42] (Methods Sec. 1.11). We focus on medication-related outcomes: hospital re-admission, all-cause mortality, and five AEs (thrombocytopenia, hyperkalemia, hypoglycemia, hyponatremia, anemia) (Methods Sec. 4.8). Using Madrigal with TransformerEHR improves performance consistently across all seven outcomes by 12.2% (AUROC) on average (Fig. 6i; Supplementary Table S19).
Oncology cohort. We curate a Dana-Farber Cancer Institute cohort of patients treated with first-line regimens comprising exactly two small-molecule oncology drugs between June 2015 and March 2025 (n=3,577; 26 two-drug regimens). We track 13 AEs across five classes: hematotoxicity, neuropathy, thromboembolism, renal impairment, and fluid/electrolyte imbalance (Supplementary Table S15; Methods Sec. 1.12).
At the population level, we quantify the association between Madrigal-predicted scores and observed AE incidence using Kendall’s $\tau$ (Methods Sec. 4.9). Madrigal scores correlate significantly with real-world incidence for 9 of 13 AEs (Kendall’s $\tau=0.26\text{-}0.68$, $p<0.05$; Supplementary Table S16). In multivariable logistic models adjusted for age, gender, race, and palliative intent, Madrigal scores remain independent predictors for 9 of 13 AEs, with positive log-odds coefficients of 0.12-1.31 (Wald $p<0.05$; Supplementary Table S17). All 13 coefficients are positive, indicating that higher Madrigal scores are consistently associated with increased AE incidence.
We then assess patient-level prediction by combining Madrigal drug embeddings with clinical covariates (age, gender, palliative intent, race, tumor tissue type) in random forest models. Performance is evaluated on a held-out test set of patients. Using Madrigal embeddings (PCA to 32 dimensions) achieves an average AUROC of 0.68 and outperforms competing methods in 9 of 13 outcomes; Madrigal outperforms Morgan fingerprints by up to 8.7% (hypercalcemia) and one-hot drug encoding approach by up to 9.8% (hypocalcemia) (Fig. 6j; Supplementary Table S18), demonstrating utility for individualized risk estimation in real-world decision-making.